Suppose a consumers demand function is dependent on both prices and income and takes the

form: Qd=500-40P+(1/10)*I (I= income here.) We also know the supply function: Qs=50P+100.

What is income elasticity when Income = 1000?
Answer #1

Demand function is Qd = 500 - 40P + (1/10)*I

At I = 1000, we have Qd = 500 - 40P + 1000/10 which gives Qd = 600 - 40P

Supply function is Qs = 50P + 100

Find equilibrium quantity

Qd = Qs

600 - 40P = 50P + 100

500 = 90P

P = 5.55

Q = 377.77

Income elasticity = dQ/dI * I/Q

= (1/10)*(1000/377.77)

= 0.265

Hence income elasticity is +0.265